The weights of bags of baby carrots are normally​distributed, with a mean of 33 ounces and a standard deviation of 0.39 ounce. Bags in the upper​4.5% are too heavy and must be repackaged. What is the most a bag of baby carrots can weigh and not need to be​repackaged?
Answer #1

solution

Using standard normal table,

P(Z > z) = 4.5%

= 1 - P(Z < z) = 0.045

= P(Z < z ) = 1 - 0.045

= P(Z < z ) = 0.955

= P(Z < 1.70 ) = 0.955  

z = 1.70 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x= 1.70 *0.39+33

x= 33.663

x=34